### v1

@@ 3.2 Sensor Configuration Optimization
Algorithm 1 Multi-LiDAR Placement Optimization
@@
7:     Update $\mathbf{m}^{(k+1)}$ based on the top $M_{k}$ best solutions $\hat{\mathbf{u}}_{i}^{(k)}$ via Eqn. (7)
8:     Update $\sigma^{(k+1)}$ and $\mathbf{C}^{(k+1)}$ via Eqn. (8), (9), (10), and (11)
@@
Objective Function. We define the objective function as $F(\mathbf{u}_{j})=\mathbf{M}_{SOG}(L_{j})$, where $\mathbf{u}_{j}\in\mathcal{U}\subset\mathbb{R}^{n}$ represents the LiDAR configuration $L_{j}$ and is subject to the physical constraint $P(\mathbf{u}_{j})=0$, and $P(\mathbf{u})>0$ if it is violated, e.g., mutual distance between LiDARs and distance from a 2D plane. Without ambiguity, we omit subscript $j$ in the following text. We then transform the constrained optimization into the unconstrained Lagrangian form as $G(\mathbf{u})=-F(\mathbf{u})+\lambda P(\mathbf{u})$, where $\lambda$ is the Lagrange multiplier. We optimize $G(\mathbf{u})$ through an iterative process that adapts the distribution of candidate solutions as Algorithm 1.
Optimization Approach. We first define a multivariate normal distribution $\mathcal{N}(\mathbf{m}^{(k)},(\sigma^{(k)})^{2}\mathbf{C}^{(k)})$, where $\mathbf{m}^{(k)}$, $\sigma^{(k)}$, and $\mathbf{C}^{(k)}$ are the mean vector, step size, and covariance matrix of the distribution of iteration $k$, respectively. We discretize the configuration space $\mathcal{U}$ with density $\delta$ and sample $N_{k}$ candidates $u_{i}^{(k)}\sim\mathcal{N}(\mathbf{m}^{(k)},(\sigma^{(k)})^{2}\mathbf{C}^{(k)})$ in each iteration $k$. We update mean vector $\mathbf{m}^{k+1}$ for the next iteration $k+1$ as the updated center of the search distribution for LiDAR configuration. The overall process can be depicted as follows:
$\mathbf{C}^{(k+1)}=(1-c_{\mathbf{C}})\mathbf{C}^{(k)}+c_{\mathbf{C}}\mathbf{p}_{\mathbf{C}}^{(k+1)}{\mathbf{p}_{\mathbf{C}}^{(k+1)}}^{T}~{},$ (9)
3.3 Theoretical Analysis
Once the evolution optimization empirically converges, it holds that the optimized solution is the local optima of the $\delta$-density Grids space of LiDAR configuration space. In this section, we provide a stronger optimality certification to theoretically ensure that the optimized placement is close to the global optimum.
Theorem 3.1 (Optimality Certification)
$\|G(\textbf{u}^{*})-G(\textbf{u}^{*}_{S})\|_{2}\leq C_{M}+k_{G}\delta~{},$ (12)
where $\textbf{u}^{*}$ is the global optima over $\mathcal{U}$.
Proof
For the $\delta$-density subset $S\subset\mathcal{U}$, it holds that $\forall\textbf{u}\in\mathcal{U}$, $\min_{\textbf{u}_{S}\in S}\|\textbf{u}-\textbf{u}_{S}\|_{2}\leq\delta$. By the Lipschitz condition, we have $\|G(\textbf{u}_{1})-G(\textbf{u}_{2})\|_{2}\leq k_{G}\|\textbf{u}_{1}-\textbf{u}_{2}\|_{2},\forall\textbf{u}_{1},\textbf{u}_{2}\in B$. By the bounded fluctuation of $G$ over $S$, it holds that $\|G(\textbf{u}_{1})-G(\textbf{u}_{2})\|_{2}\leq\|\max_{\textbf{u}\in S}G(\textbf{u})-\min_{\textbf{u}\in S}G(\textbf{u})\|_{2}\leq C_{M},\forall\textbf{u}_{1},\textbf{u}_{2}\in S$. Therefore, by absolute value inequality, we have:
[…]
where there exists a candidate $\textbf{u}_{S}$ that is subject to the condition $(\textbf{u}^{*}-\textbf{u}_{S})\leq\delta$, which concludes the proof by combining the above inequality. $\blacksquare$
The global optimality certification Theorem 3.1 is applicable in practice because the Lipschitz constant $k_{G}$ and the distance between the maximal and minimal of objective function $G$ over $S$ can be approximated easily through calculating $G(\textbf{u}_{i}^{(k)})$ of Algorithm 1 for each sampled $\textbf{u}_{i}^{(k)}$ over the $\delta$-density Grids subset. Besides, we have a more general remark below to further relax the assumption. The full proof can be found in the supplementary materials.
Remark 1
When $\mathcal{U}$ is a hyper-rectangle with the bounded $\ell_{2}$ norm of domain $U_{i}\in\mathbb{R}$ at each dimension $i$, with $i=1,2,\dots,n$, we remark Theorem 3.1 can hold in a more general way by only assuming that the Lipschitz constant $k_{G}$ of the objective function is given, where Eq. 12 becomes:
4 Experiments
@@
(g) Ours (Det)
(h) Ours (Seg)
Figure 4: Visualized LiDAR Placements. We compare the placements optimized from our proposed M-SOG metric (for both detection and segmentation tasks) and heuristic LiDAR placements utilized by major autonomous vehicle companies (cf. Fig. 1).
Table 1: Optimized M-SOG Metrics. We report $\mathbf{M}_{SOG}$ scores for both tasks. The scores are calculated based on the car class for detection and all semantic classes for segmentation. Line-roll and Pyramid-roll are abbreviated as L-roll and P-roll.
Metrics [$\mathbf{M}_{SOG}$] Center Line Pyramid Square Trapezoid L-roll P-roll Ours
@@ 4.1 Benchmark Setups
Data Generation. We generate LiDAR point clouds and the corresponding ground truth using the CARLA simulator [25]. Following previous literature [34], we use the maps of Towns 1, 3, 4, and 6 and set 6 ego-vehicle routes for each map. We incorporate 23 semantic classes for LiDAR semantic segmentation and 3 classes of objects for 3D object detection (Car, Bicycle, and Pedestrian). For each LiDAR configuration, the sub-dataset consists of 13,600 frames of samples, comprising 11,200 samples for training and 2,400 samples for validation, following the split ratio used in nuScenes [13]. More details are in the Appendix.
Corruption Types. To replicate adverse conditions, we synthesized six types of corrupted point clouds on the validation set of each sub-set, following the settings of the Robo3D benchmark [46]. These corruptions can be categorized into three groups: 1) Severe Weather Conditions, including fog, snow, and wet ground; 2) External Disturbances, such as motion blur; and 3) Internal Sensor Failure, including crosstalk and incomplete echo. See Appendix for more details.
@@
Multi-LiDAR Configurations. We adopt five commonly employed heuristic LiDAR placements, which have been adopted by several leading autonomous driving companies, as our baseline. These placements are represented in Fig. 4a-e. Following the KITTI dataset [29], we configured the LiDAR sensor with a vertical field of view (FOV) of [-24.8, 2.0] degrees. To achieve the Line-roll and Pyramid-roll configurations, we adjusted the roll angles of the LiDARs for both the Line and Pyramid setups. The Line-roll configuration is depicted in Fig. 4.
(a) Clean (Seg)
(b) Corruptions (Seg)
(c) Clean (Det)
(d) Corruptions (Det)
Figure 5: The relationship between perception performance and M-SOG, separated for the LiDAR semantic segmentation (a & b) and 3D object detection (c & d) tasks. Notably, we report true scores in clean condition and average scores under corruptions.
@@
SOG Synthesis. We follow the steps as depicted in Fig. 3 to synthesize the semantic occupancy grids for each scene. We first collect point clouds and semantic labels using one high-resolution LiDAR in CARLA and sequentially divide all samples into multiple subsets. Through the transformation of world coordinates of the ego-vehicle, frames of point clouds of each subset are aggregated into one frame of dense point cloud. Then, we utilize the voting strategy to determine the semantic label for each voxel in ROI and generate the SOG. This process is executed across all subsets to produce P-SOG. Notably, the P-SOG is only generated on the scenes of the training set.
@@
Surrogate Metric of SOG. We compute the scores of M-SOG separately for detection and segmentation, as shown in Tab. 1. To evaluate M-SOG for segmentation, we utilize all semantic classes to generate the semantic occupancy grid. For detection, we specifically focus on the car semantic type, while merging the remaining semantic types into a single category for M-SOG analysis.
[…]
Method Center Line Pyramid Square
mIoU mAcc ECE mIoU mAcc ECE mIoU mAcc ECE mIoU mAcc ECE
MinkUNet [21] $65.7$ $72.4$ $0.041$ $59.7$ $67.7$ $0.037$ $62.7$ $70.6$ $0.072$ $60.7$ $68.4$ $0.043$
PolarNet [99] $71.0$ $76.0$ $0.033$ $67.7$ $74.1$ $0.034$ $67.7$ $73.0$ $0.032$ $69.3$ $74.7$ $0.033$
SPVCNN [78] $67.1$ $74.4$ $0.034$ $59.3$ $66.7$ $0.068$ $67.6$ $74.0$ $0.037$ $63.4$ $70.2$ $0.031$
Cylinder3D [105] $72.7$ $79.2$ $0.041$ $68.9$ $76.3$ $0.045$ $68.4$ $76.0$ $0.093$ $69.9$ $76.7$ $0.044$
Average $69.1$ $75.5$ $0.037$ $63.9$ $71.2$ $0.046$ $66.6$ $73.4$ $0.059$ $65.8$ $72.5$ $0.038$
Method Trapezoid Line-Roll Pyramid-Roll Ours
mIoU mAcc ECE mIoU mAcc ECE mIoU mAcc ECE mIoU mAcc ECE
MinkUNet [21] $59.0$ $66.2$ $0.040$ $58.5$ $66.4$ $0.047$ $62.2$ $69.6$ $0.051$ $66.5$ $73.2$ $0.031$
PolarNet [99] $66.8$ $72.3$ $0.034$ $67.2$ $72.8$ $0.037$ $70.9$ $75.9$ $0.035$ $76.7$ $81.5$ $0.033$
SPVCNN [78] $61.0$ $68.8$ $0.044$ $60.6$ $68.0$ $0.034$ $67.9$ $74.2$ $0.033$ $68.3$ $74.6$ $0.034$
Cylinder3D [105] $68.5$ $75.4$ $0.057$ $69.8$ $77.0$ $0.048$ $69.3$ $77.0$ $0.048$ $73.0$ $78.9$ $0.037$
Average $63.8$ $70.7$ $0.044$ $64.0$ $71.1$ $0.042$ $67.6$ $74.2$ $0.042$ $71.1$ $77.1$ $0.034$
[…]
Method Center Line Pyramid Square
Car Ped Bicy Car Ped Bicy Car Ped Bicy Car Ped Bicy
PointPillars [50] $46.5$ $19.4$ $27.1$ $43.4$ $22.0$ $27.7$ $46.1$ $24.4$ $29.0$ $43.8$ $20.8$ $27.1$
CenterPoint [95] $55.8$ $28.7$ $28.8$ $54.0$ $34.2$ $37.7$ $55.9$ $37.4$ $35.6$ $54.0$ $35.5$ $34.1$
BEVFusion-L [58] $52.5$ $31.9$ $32.2$ $49.3$ $29.0$ $29.5$ $51.0$ $21.7$ $27.9$ $49.2$ $27.0$ $26.7$
FSTR [97] $55.3$ $27.7$ $29.3$ $51.9$ $30.2$ $33.0$ $55.7$ $29.4$ $33.8$ $52.8$ $30.3$ $31.3$
Average $52.5$ $26.9$ $29.4$ $49.7$ $28.9$ $32.0$ $52.2$ $28.2$ $31.6$ $50.0$ $28.4$ $29.8$
Method Trapezoid Line-Roll Pyramid-Roll Ours
Car Ped Bicy Car Ped Bicy Car Ped Bicy Car Ped Bicy
PointPillars [50] $43.5$ $21.5$ $27.3$ $44.6$ $21.3$ $27.0$ $46.1$ $23.6$ $27.9$ $46.8$ $24.9$ $27.2$
CenterPoint [95] $55.4$ $35.6$ $37.5$ $55.2$ $32.7$ $37.2$ $56.2$ $36.5$ $35.9$ $57.1$ $34.4$ $37.3$
BEVFusion-L [58] $50.2$ $30.0$ $31.7$ $50.8$ $29.4$ $29.5$ $50.7$ $22.7$ $28.2$ $53.0$ $28.7$ $29.5$
FSTR [97] $54.6$ $30.0$ $33.3$ $53.5$ $29.8$ $32.4$ $55.5$ $29.9$ $32.0$ $56.6$ $31.9$ $34.1$
Average $50.9$ $29.3$ $32.5$ $51.0$ $28.3$ $31.5$ $52.1$ $28.2$ $31.0$ $53.4$ $30.0$ $32.0$
4.2 Comparative Study
@@
Table 5: Robustness evaluation of 3D object detection under adverse conditions. For each LiDAR placement, we report the mAP scores for the car class, represented in percentage (%). The average scores are obtained from adverse scenarios.
@@
We conduct benchmark studies to evaluate the performance of different LiDAR placements in both clean and adverse conditions. We extensively examine the correlation between the proposed surrogate metric, known as M-SOG, and the final performance results. Through our analysis, we are able to demonstrate the effectiveness and robustness of the entire Place3D framework. Moreover, we analyze the effect of the roll angle in LiDAR placement.
@@
Effectiveness of M-SOG Surrogate Metric in Place3D. In Figs. 5(c), 5(d), 5(a) and 5(b), we examine the correlation between the proposed M-SOG surrogate metric and the perception performance in both detection and segmentation tasks. The experimental results demonstrate a clear correlation, where the performance generally improves for both tasks as the M-SOG increases. While there may be fluctuations in some placements with specific algorithms, the overall relationship follows a linear correlation, highlighting the effectiveness of M-SOG for computationally efficient sensor optimization purposes.
@@
Superiority of Optimization via Place3D. We present the object detection and semantic segmentation performance of varied LiDAR placements with several state-of-the-art algorithms in Tabs. 2, 3, 5 and 4. Notably, our optimized configurations achieved the best performance in both segmentation and detection tasks under both clean and adverse conditions among all models. The improvement from optimization remains significant even when comparing our optimized configurations against the best-performing baseline in each task. For segmentation, optimized placements outperform the best-performing baseline by up to 0.8% on clean datasets and by as much as 1.5% on corruption, highlighting the enhanced effectiveness under adverse conditions. For detection, optimized configurations exceed the best-performing baseline by up to 0.7% on clean datasets and by up to 0.3% on corruption datasets. Furthermore, we compare the performance between clean and adverse conditions. In segmentation tasks, there is a performance discrepancy of 9% on clean datasets, which goes down to 7.6% under corruptions. For detection tasks, the maximum performance gap among LiDAR placements is 3.8% on clean datasets, but this increases to 7.3% under corruptions, demonstrating an increased sensitivity to disturbances. Despite a decrease in performance under adverse conditions, the optimized LiDAR placement consistently demonstrates the best performance.
@@
Robustness of Optimization via Place3D. Under the corrupted setting, although the correlation between M-SOG and perception performance is not as obvious as that in the clean setting and shows some fluctuation, our optimized LiDAR configuration consistently maintained its performance in adverse conditions, mirroring its effectiveness in the clean condition. While the top-performing baseline LiDAR configurations in clean datasets might be notably worse compared to others when faced with corruption, the optimized configuration via Place3D consistently shows the best performance under adverse conditions.
Analysis of Roll Angle Influence. Fine-tuning the orientation angles of the LiDARs on autonomous vehicles is an effective strategy to minimize blind spots and broaden the perception range. For segmentation, Pyramid-roll slightly outperforms Pyramid, whereas Line-roll falls short of Line, which is highly aligned with scores of M-SOG in Tab. 1. For detection tasks, the situation is reversed, but still consistent with the M-SOG results. This suggests that adjusting the LiDAR’s angle can have varying impacts on the performance of object detection and semantic segmentation, necessitating a detailed analysis and optimization.

### v2

@@ 3.3 Sensor Configuration Optimization
Objective Function. We define the objective as $F(\mathbf{u}_{j})=M_{SOG}(L_{j})$, where $\mathbf{u}_{j}\in\mathcal{U}\subset\mathbb{R}^{n}$ represents the LiDAR configuration $L_{j}$ and is subject to the physical constraint $P(\mathbf{u}_{j})=0$, and $P(\mathbf{u})>0$ if it is violated, e.g., mutual distance between LiDARs and distance from a 2D plane. In our optimization, $\mathbf{u}_{j}$ includes the 3D coordinates and rolling angles of LiDARs, $\mathcal{U}$ is the finite cuboid space above the vehicle roof. Without ambiguity, we omit subscript $j$ in the following text. We then transform the constrained optimization into the unconstrained Lagrangian form as $G(\mathbf{u})=-F(\mathbf{u})+\lambda P(\mathbf{u})$, where $\lambda$ is the Lagrange multiplier. We optimize $G(\mathbf{u})$ through an iterative process that adapts the distribution of candidate solutions as Algorithm 1.
Algorithm 1 Multi-LiDAR Placement Optimization for Semantic 3D Scene Understanding
@@
7:     Update $\mathbf{m}^{(k+1)}$ based on the top $M_{k}$ best solutions $\hat{\mathbf{u}}_{i}^{(k)}$ via Equation (6)
8:     Update $\sigma^{(k+1)}$ and $\mathbf{C}^{(k+1)}$ via Equation (7), (8), (9), and (10)
@@
(g) Pyramid-roll
(h) Ours (Det)
(i) Ours (Seg)
(j) Ours (2D)
Figure 2: Visualized LiDAR Placements. We compare the LiDAR placements optimized from our proposed M-SOG metric (for LiDAR semantic segmentation and 3D object detection) and heuristic LiDAR placements utilized by major autonomous vehicle companies (see Appendix Section B.1).
Optimization Approach. We first define a multivariate normal distribution $\mathcal{N}(\mathbf{m}^{(k)},(\sigma^{(k)})^{2}\mathbf{C}^{(k)})$, where $\mathbf{m}^{(k)}$, $\sigma^{(k)}$, and $\mathbf{C}^{(k)}$ are the mean vector, step size, and covariance matrix of the distribution of iteration $k$, respectively. We discretize the configuration space $\mathcal{U}$ with density $\delta$ and sample $N_{k}$ candidates $u_{i}^{(k)}\sim\mathcal{N}(\mathbf{m}^{(k)},(\sigma^{(k)})^{2}\mathbf{C}^{(k)})$ in each iteration $k$, where $i$ indexes the candidates. We update mean vector $\mathbf{m}^{k+1}$ for the next iteration $k+1$ as the updated center of the search distribution for LiDAR configuration. The overall process can be depicted as follows:
$\mathbf{C}^{(k+1)}=(1-c_{\mathbf{C}})\mathbf{C}^{(k)}+c_{\mathbf{C}}\mathbf{p}_{\mathbf{C}}^{(k+1)}{\mathbf{p}_{\mathbf{C}}^{(k+1)}}^{T}~{}.$ (8)
3.4 Theoretical Analysis
Once the evolution optimization empirically converges, it holds that the optimized solution is the local optima of the $\delta$-density Grids space of LiDAR configuration space. In this section, we provide a stronger optimality certification to theoretically ensure that the optimized placement is close to the global optimum under the assumption of bounded and smooth objective function. Due to space limits, the full proof has been attached to the Appendix Section E.
Theorem 1 (Optimality Certification).
$\|G(\textbf{u}^{*})-G(\textbf{u}^{*}_{S})\|_{2}\leq C_{M}+k_{G}\delta~{},$ (11)
where $\textbf{u}^{*}$ is the global optima over $\mathcal{U}$.
The global optimality certification Theorem 1 is applicable in practice because the Lipschitz constant $k_{G}$ and the distance between the maximal and minimal of objective function $G$ over $S$ can be approximated easily through calculating $G(\textbf{u}_{i}^{(k)})$ of Algorithm 1 for each sampled $\textbf{u}_{i}^{(k)}$ over the $\delta$-density Grids subset. Besides, we have a more general corollary below to further relax the assumption of bounded objective function.
Corollary 1.
When $\mathcal{U}$ is a hyper-rectangle with the bounded $\ell_{2}$ norm of domain $U_{i}\in\mathbb{R}$ at each dimension $i$, with $i=1,2,\dots,n$, Theorem 1 can hold in a more general way by only assuming that the Lipschitz constant $k_{G}$ of the objective function is given, where Equation (11) becomes:
4 Experiments
Figure 3: Pipeline of Probabilistic SOG generation. We first merge multiple frames of raw point clouds (a) into dense point clouds (b). Then, we voxelize dense point clouds into SOG, i.e., semantic occupancy grids (c), and traverse all frames of dense point clouds to synthesize probabilistic SOG (d).
Table 1: Optimized M-SOG Metrics. We report the $M_{SOG}$ scores for both detection and segmentation tasks in our pipeline. The scores are calculated based on the car class for detection and all semantic classes for segmentation. Line-roll and Pyramid-roll are abbreviated as L-roll and P-roll.
Metrics [$M_{SOG}$] Center Line Pyramid Square Trapezoid L-roll P-roll Ours
@@ 4.1 Benchmark Setups
Data Generation. We generate LiDAR point clouds and ground truth using CARLA [23]. We use the maps of Towns 1, 3, 4, and 6 and set 6 ego-vehicle routes for each map. We incorporate 23 semantic classes for LiDAR semantic segmentation and 3 instance classes for 3D object detection. Data collection is performed for 10 LiDAR placements, resulting in a total of 280,000 frames: 1) For each placement, we gather 340 clean scenes and 360 corrupted scenes, with each scene consisting of 40 frames. 2) The clean set comprises 13,600 frames, including 11,200 samples (280 scenes) for training and 2,400 samples (60 scenes) for validation, following the split ratio used in nuScenes [7]. 3) The corruption set is used for robustness evaluation and contains 6 different conditions, each with 2,400 samples (60 scenes). More details are in the Appendix Section A.
@@
LiDAR Configurations. We adopt five commonly employed heuristic LiDAR placements, which have been adopted by leading autonomous driving companies, as our baseline. These placements are represented in Figure 2a-e. Following KITTI [29], we configured the LiDAR sensor with a vertical field of view of [-24.8, 2.0] degrees. To achieve the Line-roll and Pyramid-roll configurations, we adjusted LiDAR roll angles for the Line and Pyramid setups, as depicted in Figure 2. We present detailed LiDAR configurations in the Appendix Section B.1.
Corruption Types. To replicate adverse conditions, we synthesized six types of corrupted point clouds on the validation set of each sub-set, following the settings of the Robo3D benchmark [46]. These corruptions can be categorized into 1) severe weather conditions, including "snow", "fog", and "wet ground", 2) external disturbances, including "motion blur", and 3) internal sensor failures, including "crosstalk" and "incomplete echo". We include the features and implementation details of corruptions in the Appendix Section B.2.
@@
P-SOG Synthesis. We follow the steps as depicted in Figure 3 to generate semantic occupancy grids for each LiDAR scene. We first collect point clouds and semantic labels using high-resolution LiDARs and sequentially divide all samples into multiple subsets. Through the transformation of world coordinates of the ego-vehicle, frames of point clouds of each subset are aggregated into one frame of dense point cloud. Then, we utilize the voting strategy to determine the semantic label for each voxel in ROI and generate the SOG. This process is executed across all subsets to produce P-SOG. Notably, the P-SOG is only generated on the scenes of the training set.
@@
Surrogate Metric of SOG. We compute the scores of M-SOG separately for 3D detection and segmentation, as shown in Table 1. To evaluate M-SOG for segmentation, we utilize all semantic classes to generate semantic occupancy grids. For detection, we specifically focus on the car semantic type, while merging the remaining semantic classes into a single category for M-SOG analysis.
Detector and Segmentor. For the benchmark, we conduct experiments using four LiDAR semantic segmentation models, i.e., MinkUNet [18], SPVCNN [86], PolarNet [111], and Cylinder3D [117], and four 3D object detection models, i.e., PointPillars [51], CenterPoint [107], BEVFusion-L [63], and FSTR [109]. The detailed training configurations are included in the Appendix Section B.3.
@@
Table 3: Performance evaluations of 3D object detection under clean and adverse conditions. For each LiDAR placement strategy, we report the mAP scores ($\uparrow$) for the car class, represented in percentage (%). The average scores only include adverse scenarios.
@@ 4.2 Comparative Study
We conduct benchmark studies to evaluate the performance of varied LiDAR placements in both clean and adverse conditions. We extensively examine the correlation between the proposed surrogate metric, known as M-SOG, and the final performance results. Through our analysis, we are able to demonstrate the effectiveness and robustness of the entire Place3D framework.
@@
Effectiveness of M-SOG Surrogate Metric in Place3D. In Table 1, we report the scores of the M-SOG surrogate metric for various LiDAR placements. In Tables 2 and 3, we benchmark the 3D object detection and LiDAR semantic segmentation performance of varied LiDAR placements with state-of-the-art algorithms. Figure 4 illustrates the correlation between the proposed M-SOG surrogate metric and perception capacity. The results demonstrate a clear correlation, where the performance generally improves as the M-SOG increases. While there might be fluctuations in some placements with specific algorithms, the overall relationship follows a linear correlation, highlighting the effectiveness of our M-SOG for computationally efficient sensor optimization purposes. We show more plots in the Appendix Sections C.1 and C.2.
Comparisons to Existing Sensor Placement Methods. The effectiveness of LiDAR sensor placement metrics lies in the linear correlation between metrics and actual performance. We set the same ROI size and conduct a quantitative comparison with S-MIG [34] on 3D detection (as S-MIG [34] only works for 3D detection task). As shown in Figure 5, our metric exhibits better linear correlation. Since we introduce semantic occupancy information as a prior for the evaluation metric, our method more accurately characterizes the boundaries in the 3D semantic scene and addresses the issue of objects and environment occlusions when using 3D bounding boxes as priors. Moreover, our LiDAR placement method makes the first attempt to include both detection and segmentation tasks. We present additional comparisons in the Appendix Section C.3.
@@
Superiority of Optimization via Place3D. As shown in Tables 2 and 3, our optimized configurations achieved the best performance in both segmentation and detection tasks under both clean and adverse conditions among all models. We report per-class quantitative results in the Appendix Sections C.1 and C.2. The improvement from optimization remains significant even when comparing our optimized configurations against the best-performing baseline in each task. For segmentation, optimized placements outperform the best-performing baseline by up to 0.8% on clean datasets and by as much as 1.5% on corruption. For detection, optimized configurations exceed the best-performing baseline by up to 0.7% on clean datasets and by up to 0.3% on corruption.
@@
Robustness of Optimization via Place3D. The M-SOG utilizes semantic occupancy as prior knowledge, which is invariant to changing conditions, thereby enabling robustness of optimized placement in adverse weather. Under the corrupted setting, although the correlation between M-SOG and perception performance is not as obvious as that in the clean setting and shows some fluctuation, our optimized LiDAR configuration consistently maintained its performance in adverse conditions, mirroring its effectiveness in the clean condition. While the top-performing baseline LiDAR configurations in clean datasets might be notably worse compared to others when faced with corruption, the optimized configuration via Place3D consistently shows the best performance under adverse conditions. We showcase several qualitative results on this aspect in the Appendix Section D.
(a) PolarNet
(b) MinkUNet
(c) SPVCNN
(d) Cylinder3D
Figure 4: The correlation between M-SOG and LiDAR semantic segmentation [18, 86, 111, 117] models performance in the clean condition.
(a) BEVFusion (M-SOG)
(b) BEVFusion (S-MIG)
(c) PointPillars (M-SOG)
(d) PointPillars (S-MIG)
Figure 5: Comparisons of M-SOG with S-MIG [34] using BEVFusion-L [63] and PointPillars [51].
Intuitive Interpretation. We observe several intuitive reasons why refined sensor placement is beneficial in our experiments. 1) LiDAR heights. Increasing the average height of the 4 LiDARs improves performance, likely due to an expanded field of view. 2) Variation in heights. A greater difference in LiDAR heights enhances perception, as varied height distributions capture richer features from the sides of objects. 3) Uniform distribution. The pyramid placement performs better in segmentation, as a more spread-out and uniform distribution of 4 LiDARs captures a detailed surface.
(a) Against Corrupt
(b) 2D Placements (Clean)
(c) 2D Placements (Corrupt)
Figure 6: Ablation results (mIoU) of placement strategies on segmentation models [111, 18, 86, 117].